Mary sold her house for $18,000. The realtor earned 9% on the sale. If she give 3/5 of the money to her mother, keeps 2/3 for her use and puts the other 1/3 of the money in the bank. how much was the put on the bank?
Thanks

Answer by AlgebraLady88(44)

Here, we have Mary, who sold her house for $18 000. The realtor earned 9 % on the sale.
9% on the sale is 0.09 * 18 000 = $1620 (this is what the realtor earned)
So, the profit is $ 18 000 - $ 1620 = $ $16 380
Mary gives 3/5 of that to her mother.
3/5 * 16380 = $ 9828
After Mary's mother receives the $ 9828 , Mary is left with
$ 16 380 - $ 9828 = $ 6552
Mary keeps 2/3 for her use , so we find out that this amount is
2/3 * $6552 = $ 4368
She puts 1/3 into the bank and this amount should be $ 6552 - $ 4368 = $ 2184.
To check, 1/3 * $6552 = $ 2184